# Wake Vision: A Large-scale, High-Quality Dataset and Benchmark Suite for TinyML Applications

### v2

@@ Abstract
Tiny machine learning (TinyML), which enables machine learning applications on extremely low-power devices, suffers from limited size and quality of relevant datasets. To address this issue, we introduce Wake Vision, a large-scale, diverse dataset tailored for person detection—the canonical task for TinyML visual sensing. Wake Vision comprises over 6 million images, representing a hundredfold increase compared to the previous standard, and has undergone thorough quality filtering. We provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality), a smaller set with higher-quality labels. Our results demonstrate that using the Wake Vision (Quality) training set produces more accurate models than the Wake Vision (Large) training set, strongly suggesting that label quality is more important than quantity in our setting. We find use for the large training set for pre-training and knowledge distillation. To minimize label errors that can obscure true model performance, we manually label the validation and test sets, improving the test set error rate from 7.8% in the prior standard to only 2.2%. In addition to the dataset, we provide a collection of five detailed benchmark sets to facilitate the evaluation of model quality in challenging real-world scenarios that are often ignored when focusing solely on overall accuracy. These novel fine-grained benchmarks assess model performance on specific segments of the test data, such as varying lighting conditions, distances from the camera, and demographic characteristics of subjects. These benchmarks enable researchers and practitioners to better understand the strengths and limitations of their models in real-world scenarios. Our results demonstrate that using Wake Vision for training results in a 2.49% increase in accuracy compared to the established dataset. We also show the importance of dataset quality for low-capacity models and the value of dataset size for high-capacity models. The dataset, benchmark suite, code, and models are publicly available under the CC-BY 4.0 license at https://wakevision.ai/.
**footnotetext: These authors contributed equally to this work.11footnotetext: Work done while a visiting researcher at Harvard University.
@@ 1 Introduction
To reduce the energy consumption and cost of deploying Machine Learning (ML) deployments, researchers have pioneered ultra-low-power machine learning (TinyML) [3, 30, 1, 6]. TinyML co-locates ML models with the sensors by deploying the model onto always-on microcontrollers (MCUs) and TinyML accelerators, thereby reducing the energy cost of transmitting data or constantly running larger processors. Deploying models on these tiny devices imposes severe constraints, confining the model’s size to $\sim$100 kilobytes, nearly four orders of magnitude less memory than mobile phones [2].
@@
TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [5, 29] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models on these datasets.
To enable and mature the field of TinyML research, we present the Wake Vision dataset, a large and high-quality dataset for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task where a model detects whether a person is present. While simple, this task has many important use cases, such as occupancy detection [25], smart HVAC/lighting [31], or acting as an always-on ‘wake model’ in a larger ML system [11]. Wake Vision has around  6M images, ~100x more than the prior state-of-the-art person detection dataset, Visual Wake Words (VWW) [5]. Wake Vision is derived from the Open Images v7 dataset [16, 14] and is permissively licensed. Wake Vision demonstrates a 2.21% accuracy improvement over VWW.
We leverage the Wake Vision dataset for an insightful study on the trade-offs between data quality and quantity in TinyML. We connect the dataset’s unique characteristics with the limitations of TinyML models. Prior work suggests that large, overparameterized models, which have more learnable parameters than data samples, can identify and adapt to errors in their training sets [32, 9]. However, TinyML models are rarely overparameterized, leading us to hypothesize that the trade-off between data quality and quantity may be skewed towards quality for these much smaller ML models.
To explore this hypothesis and showcase the value of the Wake Vision dataset in the TinyML domain, we release two training sets as part of the Wake Vision visual corpus: Wake Vision (Large), which prioritizes dataset size, and Wake Vision (Quality), which prioritizes data quality. Using a naive training script, we find that Wake Vision (Quality) outperforms Wake Vision (Large). However, our best-performing Wake Vision model is created by using the two sets in unison, employing Wake Vision (Large) for pre-training and Wake Vision (Quality) for fine-tuning. This approach demonstrates the potential of the Wake Vision dataset to advance TinyML research by leveraging both data quantity and quality. In addition to presenting these findings, we conduct several experiments designed to investigate the trade-offs between dataset quality and quantity across a range of models. To facilitate further research into these trade-offs for TinyML, we release both training sets.
TinyML person detection systems are deployed and expected to function effectively in challenging settings, such as low lighting conditions. However, these challenging settings are often underrepresented in images sourced from the internet, which can lead to high-level metrics, like overall test set accuracy, overestimating a model’s real-world performance. To address this issue, we introduce a suite of five fine-grained benchmarks, each containing images of a specific demographic or scenario.
The benchmarks are designed to provide a more comprehensive evaluation of a person detection model’s performance across a range of real-world conditions. Our benchmarks cover several factors, including distance, which evaluates the model’s performance in detecting people at various distances from the camera; lighting, which assesses the model’s accuracy in detecting individuals in both well-lit and poorly-lit environments; depictions, which measures the model’s ability to handle different representations of people, such as drawings or sculptures; perceived gender, which investigates potential biases in the model’s performance when detecting people of different perceived genders; and age, which examines the model’s effectiveness in detecting individuals across various age groups.
By evaluating models using these fine-grained benchmarks contained within Wake Vision, we can identify potential biases, limitations, and areas for improvement that may not be apparent when relying solely on high-level metrics. This approach ensures that the person detection systems we develop are not only accurate on standard test sets but also robust and reliable across a diverse range of real-world settings and demographics. In summary, we present the following novel contributions:
  • A new person detection dataset, called Wake Vision, which is an easily accessible, permissively licensed dataset that is two orders of magnitude larger than the prior standard.
  • An improved, high-quality test set that reduces the estimated label error rate to only 2.2% from the prior standard of 7.8%
  • A suite of five fine-grained benchmark sets that evaluate a model’s fairness and robustness in challenging settings.
2 Related Work
The current de facto standard dataset for person detection is the VWW dataset [5, 6, 18, 2, 1]. To our knowledge, this is the only current open-source dataset designed explicitly for person detection. Unfortunately, the VWW dataset is small and cannot be directly downloaded. Instead, it must be regenerated from MS-COCO [19], thus limiting its accessibility and usefulness. Apart from dedicated person detection datasets, some general image classification datasets contain a person label, which makes it possible to create person detection models based on them. Examples include the Cifar-100 dataset [15], and the PASCAL Visual Object Classes dataset [8].
However, using general image classification datasets can lead to poor perceived performance of TinyML models, as they do not adequately represent the open “no-person” class [5]. Table 1 shows that our Wake Vision dataset contains up to almost two orders of magnitude more images than any other public and permissively licensed dataset. Furthermore, Wake Vision is the only person detection dataset to provide a fine-grained benchmark suite and be officially distributed to popular dataset services such as TensorFlow Datasets (TFDS) and Hugging Face Datasets. However, recent research suggests that this focus on top-line metrics can hide poor performance on critical subsets of a dataset, such as subsets relating to sensitive information and fairness [12, 13].
@@ 2 Related Work
Wake Vision (Quality) 1,322,574 624,115 9,2091 27,881
Wake Vision (Large) 5,760,428 2,880,214 - -
@@
3 Wake Vision Dataset
We introduce the Wake Vision dataset, the largest dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. We also demonstrate how its quality is significantly improved compared to existing datsets. Fig. 1 illustrates examples of images in the dataset. We discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the dataset and comparing it to the Visual Wake Words Dataset.
@@ 3.1 Label Generation
While a large person detection dataset is indispensable for TinyML research, it can be challenging to obtain. Manually labelling millions of images would be prohibitively expensive for a nascent field like TinyML. Therefore, we use existing large-scale data efforts to generate Wake Vision automatically.
The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images, which contain both image-level and bounding box labels. Image-level labels describe which objects are present in an image but not localized to a specific point in the image. Bounding box labels provide a set of four coordinates forming a box around the object. The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class. For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.”
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box. The Open Images validation and test sets are fully labeled with image-level and bounding box labels. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels). Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1).
@@
Image-level labels in Open Images contain little additional information in addition to the labels. The only additionally relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0 to 1 for labels generated by machines, and is either 0 or 1 for labels negatively or positively verified by humans respectively. See section 3.2 for more information. We use this property to exclude images from Wake Vision if they contain only low confidence person labels.
@@
Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans. This ensures that bounding box labels are less likely to be false positives. Bounding box labels can also be used to calculate an approximation of the area taken up by a person. We use this as a proxy for the distance of a person in the image, and exclude images where all person labels are far away, i.e., take up only a small portion of the image. An example of an image with a far away person can be found in Fig. 2(a). Finally, bounding box attributes include a depiction flag, which represents whether a label is a depiction. See an example of an image containing a person with the depiction flag set in Fig. 2(b). By default we use this flag to make sure that we do not consider a depiction a person.
We allow users to adapt the Wake Vision dataset to meet the needs of specific usecases by adding several configurable options in the open-source dataset creation script. For example a configuration option exists to change whether depictions are labelled as non-persons or excluded from the dataset. A comprehensive list of Open Images labels related to Wake Vision can be found in Sec. L.
@@
Training Set Comparison
To compare the performance of the two Wake Vision training sets, we train a MobileNetV2-0.25 [26] model on each dataset for an equal number of steps and compare its performance on the common test set (Table 2). The filtered training set generated from the bounding box labels outperforms the larger training set generated from the image-level labels by 4.09% test accuracy. This indicates that label quality is more important than quantity in this setting. The gap between the two is reduced to just 0.17% when training exclusively on soft labels from a teacher model (MobileNetV2-1.0) trained on bounding box labels.
Combining the two training sets to train a model where Wake Vision (Large) acts as a pre-training set and Wake Vision (Quality) acts as a fine-tuning set achieves our best performance at a mean top-level test accuracy of 85.72%. As the images in Wake Vision (Quality) is a subset of the images in Wake Vision (Large), we do not train on more images than for Wake Vision (Large).

### v3

# Wake Vision: A Tailored Dataset and Benchmark Suite for TinyML Computer Vision Applications

@@ Abstract
Tiny machine learning (TinyML) for low-power devices lacks robust datasets for development. We present Wake Vision, a large-scale dataset for person detection that contains over 6 million quality-filtered images. We provide two variants: Wake Vision (Large) and Wake Vision (Quality), leveraging the large variant for pretraining and knowledge distillation, while the higher-quality labels drive final model performance. The manually labeled validation and test sets reduce error rates from 7.8% to 2.2% compared to previous standards. In addition, we introduce five detailed benchmark sets to evaluate model performance in real-world scenarios, including varying lighting, camera distances, and demographic characteristics. Training with Wake Vision improves accuracy by 1.93% over existing datasets, demonstrating the importance of dataset quality for low-capacity models and dataset size for high-capacity models. The dataset, benchmarks, code, and models are available under the CC-BY 4.0 license, maintained by the Edge AI Foundation.
@@ 1 Introduction
Ultra-low power machine learning (TinyML) has emerged as an important technology to enable efficient ML deployments by co-locating models with sensors on microcontrollers (MCUs) [3, 38, 1, 6]. Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints. For instance, the models must often be within hundreds of kilobytes of memory in size, nearly four orders of magnitude less than mobile phones[2].
@@
TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [5, 37] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models. This creates a fundamental challenge where we can not develop robust TinyML models due to existing datasets being either too complex for tiny architectures or too limited in scale and diversity.
@@
Our work – Wake Vision (Quality) 1,322,574 624,115 9,291 27,881
Our work – Wake Vision (Large) 5,760,428 2,880,214 -- --
@@
To address this critical gap, we present Wake Vision, a large-scale dataset specifically designed for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task that, while simple, enables crucial applications ranging from occupancy detection [31] and smart HVAC/lighting [39] to acting as an always-on ’wake model’ in larger ML systems [15].
Wake Vision advances the state-of-the-art by providing 6M images, approximately 100x more than the current leading dataset Visual Wake Words (VWW) [5], while achieving a 1.93% accuracy improvement. The dataset is derived from Open Images v7 [21, 19] and is released under a permissive CC-BY 4.0 license to facilitate broad adoption.
A key contribution of our work is the first systematic investigation of how dataset quality versus quantity impacts TinyML model performance. This question is particularly relevant for resource-constrained models, as prior work suggests that large, overparameterized models can effectively identify and adapt to training set errors [40, 9]. TinyML models, with their limited capacity, may exhibit fundamentally different behavior. To explore this hypothesis, we introduce two complementary training sets within the Wake Vision visual corpus: Wake Vision (Large), emphasizing scale, and Wake Vision (Quality), prioritizing label accuracy. Our experiments reveal that while Wake Vision (Quality) outperforms its larger counterpart in naive training scenarios, optimal performance is achieved through a training strategy that leverages both sets — using Wake Vision (Large) for pre-training followed by using Wake Vision (Quality) fine-tuning. To further explore the dataset quality and quantity trade-off for tinyML, we train a range of ResNet [14] style models across Wake Vision datasets of varying error rates.
Our experiments collectively reveal that small models are more sensitive to higher error rates than large models and that large models can benefit from larger datasets despite a higher error rate. These findings provide insight into the unique challenges of training highly constrained models.
Beyond raw performance metrics, real-world deployment of TinyML systems often demands robustness across diverse and challenging use case conditions. Internet-sourced datasets often underrepresent challenging scenarios, such as low-lighting conditions, leading to overly optimistic performance estimates based on standard test set metrics.
To enable a rigorous evaluation of real-world capabilities, we additionally develop a comprehensive suite of five fine-grained benchmarks based on our Wake Vision dataset. Our benchmarks systematically evaluate performance across multiple dimensions: distance (detection at varying ranges), lighting (performance in well-lit and poorly-lit environments), depictions (handling of drawings and sculptures), perceived gender (investigating potential gender biases), and perceived age (examining effectiveness across age groups).
This benchmarking framework addresses a gap in current TinyML assessment methods and enables researchers to identify potential biases, limitations, and areas for improvement that might be masked by aggregate metrics alone. We find that Wake Vision models outperform VWW on 13 of the 16 fine-grained sets, demonstrating robust performance further illustrating the advantages of a large diverse training dataset.
In summary, we present the following contributions:
  • A production grade computer vision dataset for TinyML containing over 6M images (100$\times$ larger than state-of-the-art) for person detection, with comprehensive filtering, permissive licensing, and accessibility through standard ML frameworks.
  • A cleaned test set with 2.2% label error rate, enabling more reliable evaluation of TinyML models.
  • A benchmark suite for systematically evaluating model robustness across five critical dimensions (perceived gender, age, distance, lighting, depictions), exposing potential failure modes in real-world deployments.
  • Evidence that TinyML models are more sensitive to dataset quality than larger models, providing guidance for dataset curation in resource-constrained settings.
  • A complete, maintained dataset infrastructure including benchmarks, evaluation code, and reproducible filtering pipeline to support ongoing TinyML research.
@@
2 Background and Related Work
Person Detection and Its Role in TinyML
Person detection is a fundamental computer vision task whose job is to determine the presence or absence of humans in visual data. Unlike more complex tasks such as pose estimation or instance segmentation, person detection focuses on the binary classification problem.
This seemingly simple task has emerged as a canonical use case for TinyML systems [3, 1], driven by both theoretical foundations and practical considerations. In the context of resource-constrained computing, TinyML systems demand tasks that carefully balance computational feasibility with real-world utility. Person detection represents an optimal compromise. It is sufficiently lightweight to be implemented on embedded microcontrollers while maintaining enough discriminative power for practical applications.
The significance of person detection extends beyond its computational characteristics, serving as a critical selective activation mechanism in resource-constrained systems. These lightweight models efficiently manage the activation of power-intensive sensors and processors by determining human presence. This architectural pattern has demonstrated particular value in diverse applications, including occupancy detection [31], smart HVAC systems [39], and as an always-on "wake model" in larger ML pipelines [15]. Moreover, this approach enables substantial energy savings while preserving privacy through on-device processing.
Existing Person Detection Datasets
In the TinyML computer vision domain, the VWW dataset [5, 6, 24, 2, 1] has established itself as the de facto standard. It remains the only open-source dataset specifically designed for person detection with direct commercial applicability and potential for real-world impact. However, VWW faces significant limitations: its small size and indirect access requirements (requiring regeneration from MS-COCO [25]) restrict its accessibility and utility.
Although other data sets contain person labels within general image classification collections, such as Cifar-100 [20] and PASCAL Visual Object Classes [8], these present significant drawbacks for TinyML applications. Specifically, their inadequate representation of the open “no-person” class can lead to poor perceived performance of TinyML models [5]. Our Wake Vision dataset addresses these limitations by providing nearly two orders of magnitude more images than any existing public, permissively licensed dataset (Table 1). Furthermore, it distinguishes itself as the only person detection dataset offering a fine-grained benchmark suite and official distribution through popular services like TensorFlow Datasets (TFDS) and Hugging Face Datasets.
3 Wake Vision: Dataset and Benchmark Suite
We introduce Wake Vision, a new dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is almost two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. Fig. 1 illustrates examples of images in the dataset. In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the data set and comparing it with VWW Dataset.
@@ 3.1 Label Generation
A large person detection dataset is indispensable for TinyML research. It can, however, be challenging to obtain. Manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML. Therefore, we bootstrap Wake Vision using existing large-scale data efforts.
The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images [21, 19], which contain both image-level and bounding box labels of various objects. Image-level labels describe which objects are present in an image but not localized to a specific point in the image. Bounding box labels provide a set of four coordinates forming a box around the object.
The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class. For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.” A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix L.
Users can adapt Wake Vision to meet the needs of specific use cases by adding several configurable options to the open-source dataset creation script. For example, a configuration option exists to change whether depictions are labeled as non-persons or excluded from the dataset. We refer to Appendix I for details about these configuration options.
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels). Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1). The Open Images validation and test sets are fully labeled with bounding box labels. Therefore, both the Wake Vision validation and test sets are created using the higher-quality bounding box labels.
@@
Image-level labels in Open Images contain little additional information than the raw labels. The only additional relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0 to 1 for labels generated by machines and is either 0 or 1 for labels negatively or positively verified by humans, respectively. See Sec. 3.2 for more information. We use this property to exclude images from Wake Vision if they contain only low-confidence person labels.
@@
Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans. This ensures that bounding box labels are less likely to be false positives. Bounding box labels can be used to calculate an approximation of the area taken up by a person. We use this as a proxy for the distance of a person in the image, and by default exclude images where all person labels are far away, i.e., take up less than 5% of the image. An example of an image with a faraway person can be found in Fig. 2(a). Finally, bounding box attributes include a depiction flag, which represents whether a label is a depiction. See an example of an image containing a person with the depiction flag set in Fig. 2(b). By default, we use this flag to make sure that we do not consider a depiction a person. See Appendix C for an illustration of the bounding box filtering process.
@@
3.2 Label Correction
Label errors are a challenge in computer vision that limit progress and obscure true performance [30, 4]. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision. All labels in the Open Images dataset originate from machine-generated candidate labels. Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases. Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present but lacks the corresponding label. Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking. Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset can be used to accurately evaluate a model’s performance.
We use the Scale Rapid tool from Scale AI (https://scale.com/) to relabel the Wake Vision validation and test sets. Details about the relabelling process, including instructions given and total cost, can be found in Appendix N.
With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [29]. We perform an initial exploration of these methods in appendix O.
We report our estimated label error rate after label corrections versus VWW in Table 2. These estimates are all based on a random subset of 500 samples from each dataset, which the authors manually relabelled according to whether we saw a non-depiction person or body part in the image. The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
Table 2: Label error rate of VWW and Wake Vision after label correction. The Wake Vision error rate is broken down by set. The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different label sources. The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.
Dataset Wake Vision Set Label Error
Visual Wake Words Train, Val, & Test 7.8%
Wake Vision Train (Large) 15.2%
Wake Vision Train (Quality) 6.8%
Wake Vision Val & Test 2.2%
3.3 Fine-grained Benchmark Suite
While Wake Vision provides a substantial improvement in scale and quality over existing datasets, we recognize that overall test set performance alone may not capture the nuanced challenges of real-world deployment for TinyML applications. Internet-sourced images typically exhibit a distinct bias towards well-composed photographs—they are often well-lit, properly framed, and feature clearly visible subjects. This stands in stark contrast to the challenging conditions where TinyML person detection systems are typically deployed, such as varying lighting conditions, extreme viewing distances, or partial occlusions [32].
To bridge this gap between benchmark performance and practical utility, we augment Wake Vision with a comprehensive fine-grained benchmark suite. This suite evaluates model robustness across challenging real-world scenarios where traditional accuracy metrics might mask significant failure modes. The benchmarks assess performance across critical dimensions, including lighting conditions, subject distance, and demographic attributes, enabling developers to identify potential biases or limitations during the design phase rather than after deployment. By providing these targeted evaluation sets, we enable a more nuanced analysis of model behavior beyond aggregate accuracy metrics.
The suite consists of five fine-grained benchmark sets. Each of the five fine-grained benchmarks has been chosen based on a combination of its relevance to TinyML use cases and the availability of requisite metadata to generate the sets. Each benchmark set is a subset of the respective validation or test set filtered based on the criteria under test. The benchmark determines whether a model is sufficiently accurate in the planned deployment setting. For example, a model designer may make different design choices for a use case where the subject is close to the camera and well lit compared to the inverse setting. Appendix D contains a case study that shows how fine-grained evaluation sets can be used to design robust models.
Perceived Gender and Age
Table 3: Accuracy on the Wake Vision (WV) and VWW test sets by models trained on the VWW, Wake Vision (Quality) and Wake Vision (Combined) training sets.
Train
VWW Wake Vision (Quality) WV (Combined)
Test VWW 88.33$\pm$0.2988.59$\pm$0.1789.34$\pm$0.02%
WV 83.79$\pm$0.2384.89$\pm$0.1185.72$\pm$0.04%
Underrepresented sub-groups typically constitute a small portion of a generic test set; therefore, top-line metrics often obscure a bias in a model until it is deployed [16]. This benchmark, generated from the Open Images More Inclusive Annotations for People (MIAP) extended fairness labels [34], evaluates a model separately on demographics that are underrepresented in the underlying dataset distribution, identifying bias. These labels are based on perceived gender and age representation and are not necessarily representative.
Distance
This tests how the distance of people in images impacts model performance. If a person detection system is intended to recognize subjects at great distances, the system’s performance on the faraway dataset will be more informative than its performance on the top-level test set. We created three data sets based on the percentage of the image the subject bounding box covers. The three sets are near (>60%), at a medium distance (10-60%), and far away (<10%).
Lighting
The lighting data sets test the performance of Machine Learning (ML) systems in different lighting conditions. In scenarios like security monitoring, outdoor robotics, or augmented reality applications, models must be robust to varying lighting conditions, including low-light environments. We create three fine-grained datasets of this type for dark, normal, and bright lighting conditions, respectively. We quantify lighting conditions by the average pixel values of images in greyscale, a simple but effective method for distinguishing lighting conditions [41]. We define low as an average pixel value less than 85, normal as between 85 and 170, and bright lighting conditions as greater than 170.
Depictions
A particularly challenging task for a person detection model is to correctly reject depictions of people. In many use cases, a person detection model should not trigger on a depiction. For example, a room occupancy detector could incorrectly identify a painting on the wall as a person. This benchmark measures a model’s accuracy on three related sets of non-person samples: depictions of people, depictions that are not of people, and images that do not contain a depiction of any kind. Depictions of people can range from photo-realistic to crude stick figures.
3.4 Usability
Wake Vision and its fine-grained benchmark suite are available through TensorFlow Datasets [36] and HuggingFace Datasets [22] to enable easy access and use by the community. The images and labels are rehosted to ensure the dataset will not be changed over time due to dead links, making it a more stable benchmark. The rehosted labels are generated according to our default dataset configuration, further described in Appendix I.
4 Results
In this section, we present a comprehensive evaluation of Wake Vision through five complementary analyses. First, we provide the first insights into best practices for Wake Vision by comparing its two training sets, which also provide initial insights into the trade-off between dataset quality and size in a TinyML setting. Second, we show that Wake Vision serves as an effective drop-in replacement for Visual Wake Words by conducting cross-evaluation studies using identical MobileNetV2 models. Third, we validate Wake Vision’s benefits across diverse model architectures commonly used in the TinyML community, from MCUNet to ColabNAS families. Fourth, we thoroughly investigate the critical trade-off between dataset quality and size for TinyML models, providing insights into how model capacity influences sensitivity to label noise versus dataset scale. Finally, we present fine-grained benchmark results across demographics, environmental conditions, and visual contexts to assess model robustness in real-world scenarios. Throughout these analyses, we demonstrate that Wake Vision not only improves upon VWW’s performance but also enables a more thorough evaluation of TinyML computer vision models.
4.1 Training Set Evaluation